OpenStudy (anonymous):

HELP!!! MEDAL TO WHOEVER ANSWERS THIS QUESTION!!! What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (anonymous):

Use elimination or substitution. Though elimination would be easier.

OpenStudy (anonymous):

(3x − 6y = 3)(times)7 (7x − 5y = −11)(times)-3 ~~~~~~~~~~~~~~~~ 21x - 42y = 21 -21x + 15y = 33 Add. -27y = 54 Divide y=-2 ~~~~~~~~~~~~~~~ Substitute in y 3x − 6(-2) = 3 3x + 12 = 3 Subtract 3x = -9 Divide x = -3

OpenStudy (anonymous):

THANK YOU SOOO MUCH!!! @RoseDryer can u pls help me with 1 more question??

OpenStudy (anonymous):

Yepp.

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

@patty_emily just follow the method that @RoseDryer used in the previous question. You know a user can't just give out the answer(s).

OpenStudy (anonymous):

I know its just i still dont get it and i really need to do this @some_someone

OpenStudy (anonymous):

well if you don't get it; you can ask question(s). U can't just ask for the answer.

OpenStudy (anonymous):

(2x − y = 6)(times)-2 -4x +2y = -12 4x − 2y = 12 If we add there would be no solution because everything cancels out. Well the reason I multiply the equation is to get the two equations to cancel out a variable. So when you have 2x + 2y = 4 and 2x -4y = 8 To get rid of y and solve for x you would first multiply the first equation by 2 to get 4x +4y = 16 if you add the two equations together 4x + 4y = 16 2x - 4y = 8 you get 6x = 24 Because + 4y + (-4y) = 0

OpenStudy (anonymous):

Do you understand better now?

OpenStudy (anonymous):

Here this might help too http://www.mathsteacher.com.au/year10/ch04_simultaneous/03_elimination_method/elim.htm It shows/ explains how to do it.